OpenStudy (anonymous):

What is the solution of y − 4x = 0 and 3x + 6y = 9? x=0 , y=3 2 x=1 3 , y=4 3 x=1 , y=−1 x=1 4 , y=2 3

OpenStudy (gorv):

x=1/3 and y=4/3

OpenStudy (gorv):

@keketooskiny